An insulated box containing helium gas falls from a balloon 4.5 km above the earth’s surface. Calculate the temperature rise in oC of the helium when box hits the ground.

Temperature rise = ____ oC


1
Expert's answer
2021-09-12T16:26:35-0400

The law of transformation of energy says: decreasing in mechanical energy = increasing of internal energy. So

"mgh=\frac{3}{2}\frac{m}{M}R\Delta T"

"\Delta T=\frac{2}{3}\frac{gMh}{R}=\frac{2}{3}\frac{9.8*0.004*4500}{8.3}=14^{\circ}\rm C"